# Hyperphantasia: A Benchmark for Evaluating the Mental Visualization Capabilities of Multimodal LLMs

### v1

@@ Abstract
Mental visualization, the ability to construct and manipulate visual representations internally, is a core component of human cognition and plays a vital role in tasks involving reasoning, prediction, and abstraction. Despite the rapid progress of Multimodal Large Language Models (MLLMs), current benchmarks primarily assess passive visual perception, offering limited insight into the more active capability of internally constructing visual patterns to support problem solving. Yet mental visualization is a critical cognitive skill in humans, supporting abilities such as spatial navigation, predicting physical trajectories, and solving complex visual problems through imaginative simulation. To bridge this gap, we introduce Hyperphantasia, a synthetic benchmark designed to evaluate the mental visualization abilities of MLLMs through four carefully constructed puzzles. Each task is procedurally generated and presented at three difficulty levels, enabling controlled analysis of model performance across increasing complexity. Our comprehensive evaluation of state-of-the-art models reveals a substantial gap between the performance of humans and MLLMs. Additionally, we explore the potential of reinforcement learning to improve visual simulation capabilities. Our findings suggest that while some models exhibit partial competence in recognizing visual patterns, robust mental visualization remains an open challenge for current MLLMs. Our dataset is publicly available at Huggingface111https://huggingface.co/datasets/shahab7899/Hyperphantasia, and the evaluation code can be found at GitHub222https://github.com/AIF4S/Hyperphantasia.
@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated strong performance on a wide range of linguistic and reasoning tasks, mostly driven by scale and training on massive text corpora. Building on this success, vision-language models (VLMs) – which fuse textual and visual representations – have extended these capabilities to multimodal domains. VLMs have achieved impressive results on benchmarks that assess a variety of skills, including visual problem-solving [33], visual Math reasoning [19], and visual question answering (VQA) [9]. These benchmarks have become de facto standards for evaluating VLM performance. However, the majority of these tasks involve directly using the information given in the image, offering limited insight into the models’ ability to simulate, extrapolate, or reason about the latent structure of a scene.
@@ 3.1 Overview of Hyperphantasia
The tasks in Hyperphantasia are organized into two main categories: Interpolation and Extrapolation, each comprised of two sub-categories. The Interpolation tasks evaluate a model’s ability to infer internal boundaries and recognize visual concepts from partial information. The Extrapolation tasks assess the model’s capacity to anticipate and extend visual structures beyond the information explicitly provided. In the following sections, we describe each category along with the corresponding puzzles. Figure 1 illustrates examples of each puzzle in our benchmark.
@@ 4.1 Experimental Setup
Models: We evaluate Hyperphantasia on a range of state-of-the-art MLLMs, including o4-mini, GPT-4o, Gemini 2.5 Pro Preview, Claude 3.7 Sonnet, Qwen-VL-2.5 (7B and 32B) [2], LLaMA 3.2 (11B and 90B) [10], LLaVA-OneVision (7B and 72B) [16], Molmo (7B and 72B) [5], and Deepseek-VL2 [30]. To evaluate large open-source models (above 11B), we use four NVIDIA H100 GPUs; for Deepseek-VL 2, we use three NVIDIA A100 GPUs; and for the remaining open-source models, we use a single A100 GPU. For proprietary models, we use their APIs, which do not require any significant compute resources. Evaluation time ranges from 1 to 3 hours, depending on the model.
Human Evaluation: For Seven Segments, each digit is clearly recognizable based on its edges, making the task trivial for humans. We manually evaluated several Seven Segments puzzles and found them consistently easy to solve, so we report accuracy on this task. For the remaining puzzles, we recruited human participants, ensuring that each question was answered by at least three individuals. We report the average human accuracy across all tasks and difficulty levels. For Connect the Dots problems, since all difficulty variants are derived from the same base examples, we assigned different participants to different difficulty levels to avoid bias.
@@ 4.2 Results
Most strikingly, all models fail on Seven Segments puzzles except for Gemini and o4-mini, which achieve and accuracy on the easy set, respectively. However, the rest of the models cannot even solve one puzzle, regardless of difficulty. Upon inspecting model responses, we observe that some models repeatedly output fixed guesses regardless of the input. For example, Qwen-VL 2.5 7B answered "" in out of Easy examples. This behavior suggests that these models are not engaging with the underlying visual reasoning task at all and instead resort to memorized patterns or default completions, further underscoring their lack of mental visualization ability. Examples of this behavior are provided in the Appendix 7.2.
@@
Unlike Seven Segments puzzles, models perform relatively well on Connect the Dots puzzles in the Easy setting, with Gemini and GPT-4o reaching . However, performance degrades significantly on Medium and Hard, with Gemini being the best model with accuracy. While human accuracy also declines slightly with increased difficulty, the drop is small (around ) and the accuracy consistently remains above . Furthermore, despite the relative simplicity of the task, models such as Molmo and Deepseek struggle with Medium and Hard sets and perform near the level of random guessing. Upon closer inspection of model outputs, we find that some models confidently justify incorrect answers by hallucinating features related to the wrong option, which are not present in the image. This suggests that, rather than grounding their answers in the visual input, models often rely on loosely associated or imagined patterns, revealing limitations in both visual grounding and fine-grained perceptual reasoning. We provide examples of such answers in the Appendix 7.2.
@@
Extrapolation tasks also prove to pose a major challenge to the models. In Linear Trajectory puzzles, Claude achieves in the Easy set, the highest among all models, but drops to the level of random guessing on Medium and Hard. o4-mini shows a similar trend, scoring on Easy but struggling with more difficult examples. Interestingly, LLaVA-OneVision 72B, which performs modestly on Connect the Dots puzzles, performs on par with 4o-mini on Linear Trajectory across all difficulty levels. Overall, except for Claude, all models have below accuracy in Easy Linear Trajectory puzzles. This is particularly notable given that the human accuracy on Easy problems is . In this puzzle, we see a larger drop in human accuracy as the difficulty level increases, but the accuracy remains fairly high ( on Hard).
@@
Parabolic Trajectory puzzles are even more difficult for the models. Although Claude again has the best performance in the Easy set, its accuracy rapidly deteriorates in the Medium and Hard sets. Moreover, almost all models fall to near random guessing on the Medium and Hard sets. One exception is LLaMA 3.2 72B, which achieves in Medium Parabolic Trajectory despite underperforming on earlier tasks. LLaVA-OneVision continues to perform competitively, again on par with o4-mini and outperforming the remaining models. A noteworthy trend we observe in responses to extrapolation tasks is that some models resort to overly simplistic and incorrect heuristics, such as selecting the rightmost point as the impact location. This further demonstrates a failure to perform genuine trajectory reasoning. Examples of this behavior are provided in Appendix 7.2. Notably, human accuracy on Easy examples is close to , but drops substantially to around on Medium and Hard. We hypothesize that participants can eliminate two implausible options but often struggle to confidently choose between the final two, leading to a near-random guessing performance.
@@
Overall, these results highlight the limitations of current state-of-the-art MLLMs in performing mental visualization tasks. Even the best-performing models struggle with tasks that require mentally constructing or extending visual structures, which are intuitive and nearly trivial for humans. This persistent gap emphasizes the need for deeper investigation into how MLLMs can acquire the ability to reason over visual abstractions.
@@ Acknowledgements
We would like to thank Microsoft for an Accelerating Foundation Models Research grant that provided the OpenAI credits enabling this work. This research is also in part supported by AWS credits through an Amazon Faculty research award and a NAIRR Pilot award. M. Soltanolkotabi is also supported by the Packard Fellowship in Science and Engineering, a Sloan Research Fellowship in Mathematics, an NSF-CAREER under award #1846369, DARPA FastNICS program, and NSF-CIF awards #1813877 and #2008443. and NIH DP2LM014564-01.

### v2

@@ Abstract
Mental visualization, the ability to construct and manipulate visual representations internally, is a core component of human cognition and plays a vital role in tasks involving reasoning, prediction, and abstraction. Despite the rapid progress of Multimodal Large Language Models (MLLMs), current benchmarks primarily assess passive visual perception, offering limited insight into the more active capability of internally constructing visual patterns to support problem solving. Yet mental visualization is a critical cognitive skill in humans, supporting abilities such as spatial navigation, predicting physical trajectories, and solving complex visual problems through imaginative simulation. To bridge this gap, we introduce Hyperphantasia, a synthetic benchmark designed to evaluate the mental visualization abilities of MLLMs through four carefully constructed puzzles. Each puzzle is procedurally generated and presented at three difficulty levels, enabling controlled analysis of model performance across increasing complexity. Our comprehensive evaluation of state-of-the-art models reveals a substantial gap between the performance of humans and MLLMs. Additionally, we explore the potential of reinforcement learning to improve visual simulation capabilities. Our findings suggest that while some models exhibit partial competence in recognizing visual patterns, robust mental visualization remains an open challenge for current MLLMs. Our dataset is publicly available at Huggingface111https://huggingface.co/datasets/shahab7899/Hyperphantasia, and the evaluation code can be found at GitHub222https://github.com/AIF4S/Hyperphantasia.
@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated strong performance on a wide range of linguistic and reasoning tasks, mostly driven by scale and training on massive text corpora. Building on this success, vision-language models (VLMs), which fuse textual and visual representations, have extended these capabilities to multimodal domains. VLMs have achieved impressive results on benchmarks that assess a variety of skills, including visual problem-solving [33], visual Math reasoning [19], and visual question answering (VQA) [9]. These benchmarks have become de facto standards for evaluating VLM performance. However, the majority of these tasks involve directly using the information given in the image, offering limited insight into the models’ ability to simulate, extrapolate, or reason about the latent structure of a scene.
@@ 3.1 Overview of Hyperphantasia
The tasks in Hyperphantasia are organized into two main categories: Interpolation and Extrapolation, each comprising of two sub-categories. The Interpolation tasks evaluate a model’s ability to infer internal boundaries and recognize visual concepts from partial information. The Extrapolation tasks assess the model’s capacity to anticipate and extend visual structures beyond the information explicitly provided. In the following sections, we describe each category along with the corresponding puzzles. Figure 1 illustrates examples of each puzzle in our benchmark.
@@ 4.1 Experimental Setup
Models: We evaluate Hyperphantasia on a range of state-of-the-art MLLMs, including o4-mini (2025-04-16), GPT-4o (2024-08-06), Gemini 2.5 Pro (June 2025 update), Claude 3.7 Sonnet (20250219), Qwen-VL-2.5 (7B and 32B) [2], LLaMA 3.2 (11B and 90B) [10], LLaVA-OneVision (7B and 72B) [16], Molmo (7B and 72B) [5], and Deepseek-VL2 [30]. To evaluate large open-source models (above 11B), we use four NVIDIA H100 GPUs; for Deepseek-VL 2, we use three NVIDIA A100 GPUs; and for the remaining open-source models, we use a single A100 GPU. For proprietary models, we use their APIs, which do not require any significant compute resources. Evaluation time ranges from 1 to 3 hours, depending on the model.
Human Evaluation: For Seven Segments, each digit is clearly recognizable based on its edges, making the task trivial for humans. Moreover, we manually evaluated several Seven Segments puzzles and found them consistently easy to solve, so we report accuracy on this task. For the remaining puzzles, we recruited human participants while ensuring that each question was answered by at least three individuals. We report the average human accuracy across all tasks and difficulty levels. For Connect the Dots problems, since all difficulty variants are derived from the same base examples, we assigned different participants to different difficulty levels to avoid biases.
@@ 4.2 Results
Most strikingly, all models fail on Seven Segments puzzles except for Gemini and o4-mini, which achieve and accuracy on the easy set, respectively. However, the rest of the models almost completely fail, regardless of difficulty. Upon inspecting model responses, we observe that some models repeatedly output fixed guesses regardless of the input. For example, Qwen-VL 2.5 7B answered "" in out of Easy examples. This behavior suggests that these models are not engaging with the underlying visual reasoning task at all and instead resort to memorized patterns or default completions, further underscoring their lack of mental visualization ability. Examples of this behavior are provided in the Appendix 8.
@@
Unlike Seven Segments puzzles, models perform relatively well on Connect the Dots puzzles in the Easy setting, with Gemini and GPT-4o reaching and , respectively. However, performance degrades significantly on Medium and Hard, with Gemini being the best model with accuracy on the Hard set. While human accuracy also declines with increased difficulty, the drop is much smaller (around ) and the accuracy consistently remains above . Furthermore, despite the relative simplicity of the task, models such as Molmo and Deepseek struggle with Medium and Hard sets and perform near the level of random guessing. Upon closer inspection of model outputs, we find that some models confidently justify incorrect answers by hallucinating features related to the wrong option, which are not present in the image. This suggests that, rather than grounding their answers in the visual input, models often rely on loosely associated or imagined patterns, revealing limitations in both visual grounding and fine-grained perceptual reasoning. We provide examples of such answers in the Appendix 8.
@@
Extrapolation tasks also prove to pose a major challenge to the models. In Linear Trajectory puzzles, Claude achieves in the Easy set, the highest among all models, but drops to the level of random guessing on Medium and Hard. o4-mini shows a similar trend, scoring on Easy but struggling with more difficult examples. Interestingly, LLaVA-OneVision 72B, which performs modestly on Connect the Dots puzzles, performs on par with 4o-mini on Linear Trajectory Easy difficulty level. Overall, except for Claude, all models have below accuracy in Easy Linear Trajectory puzzles. The struggle with this task is particularly notable given that the human accuracy on Easy problems is . In this puzzle, we see a larger drop in human accuracy as the difficulty level increases, but the accuracy remains fairly high ( on Hard).
@@
Parabolic Trajectory puzzles are even more difficult for the models. Although Claude again has the best performance in the Easy set, its accuracy rapidly deteriorates in the Medium and Hard sets. Moreover, almost all models fall to near random guessing on the Medium and Hard sets. One exception is LLaMA 3.2 72B, which achieves in Medium Parabolic Trajectory despite underperforming on earlier tasks. LLaVA-OneVision continues to perform competitively, again on par with o4-mini and outperforming the remaining models. A noteworthy trend we observe in responses to extrapolation tasks is that some models resort to overly simplistic and incorrect heuristics, such as selecting the rightmost point as the impact location. This further demonstrates a failure to perform genuine trajectory reasoning. Examples of this behavior are provided in Appendix 8. Notably, human accuracy on Easy examples is close to , but drops substantially to around on Medium and Hard. We hypothesize that participants can eliminate two implausible options but often struggle to confidently choose between the final two, leading to a near-random guessing performance.
@@
Overall, these findings underscore the limitations of current state-of-the-art MLLMs in performing mental visualization tasks, revealing a significant gap between models and human capabilities. Even the best-performing models struggle with tasks that require mentally constructing or extending visual structures, which are intuitive and nearly trivial for humans. This persistent gap emphasizes the need for deeper investigation into how MLLMs can acquire the ability to reason over visual abstractions.
@@ 5.3 Visual Recognition Failures in Seven Segments
5.4 Future Directions
Beyond reinforcement techniques, we believe that a promising direction is to equip models with visual thinking capabilities. Current models reason entirely in the language domain, but Hyperphantasia puzzles and many real-world scenarios, the thinking cannot be explained with language, and they require visual thinking. While recent "omni" models offer some potential in this area, they are still significantly behind state-of-the-art language models in overall performance.
Importantly, we do not believe that omni models are the only path forward. Visual thinking tokens do not need to correspond to meaningful or interpretable images. Instead, they could function as internal visual representations. However, designing and training such models requires careful consideration, especially in how to cue or supervise this form of internal visualization. Some early efforts in this direction, such as [1], have explored visual reasoning in constrained setups, but these remain narrow in scope. We see Hyperphantasia as an ideal testbed to encourage progress in this direction.
@@ Acknowledgements
We would like to thank Microsoft for an Accelerating Foundation Models Research grant that provided the OpenAI credits enabling this work. This research was also in part supported by AWS credits through an Amazon Faculty Research Award and a NAIRR Pilot Award. M. Soltanolkotabi and MS. Sepehri were supported by the USC–Capital One Center for Responsible AI and Decision Making in Finance (CREDIF) Fellowship. M. Soltanolkotabi is also supported by the Packard Fellowship in Science and Engineering, a Sloan Research Fellowship in Mathematics, an NSF-CAREER under award #1846369, DARPA FastNICS program, and NSF-CIF awards #1813877 and #2008443. and NIH DP2LM014564-01.
Appendix74
7 Evaluation Prompts
In this part, we provide the exact prompts used for the evaluation of each task.
7.1 Seven Segments
For these puzzles, the edges to connect and the number of digits vary among puzzles, but the rest of the prompt is the same for all samples.
Solve the following puzzle. The rules are as follows.
1. You see a grid of dots.
2. Every dot is denoted by a red number on top right of it.
3. Connect the following dots:
**EDGES TO CONNECT**
4. What is the **N**-digit number formed in the image after drawing these connections?
You can explain your reasoning but you should specify your final answer by putting <ANSWER> before and after it like <ANSWER>FINAL_ANSWER<ANSWER> where FINAL_ANSWER is your
answer. For example, if your answer is A, you should say <ANSWER>A<ANSWER>
7.2 Connect the Dots
In these prompts, the options vary for different puzzles.
In the image you see a set of dots. These dots were sampled from the outer contour of an image. Based on these dots, which of these options best describes the image?
A: **OPTION A**
B: **OPTION B**
C: **OPTION C**
D: **OPTION D**
Answer with the letter of the option that you think is correct.
You can explain your reasoning but you should specify your final answer by putting <ANSWER> before and after it like <ANSWER>FINAL_ANSWER<ANSWER> where FINAL_ANSWER is your
answer. For example, if your answer is A, you should say <ANSWER>A<ANSWER>
7.3 Linear Trajectory
We use the same prompt for all of the puzzles.
The image shows a triangle composed of two blue edges and one red edge. A portion of the red edge has been removed. The red line originates from the left red vertex and should connect to a point along the blue edge on the right side of the triangle. There are four possible points marked on this blue edge: A, B, C, and D. Question: Based on the direction and alignment of the red edge, extrapolate the rest of the red line and determine which point (A, B, C, or D) is the vertex of the triangle corresponding to the missing part of the red edge? Select the most geometrically consistent option.
You can explain your reasoning but you should specify your final answer by putting <ANSWER> before and after it like <ANSWER>FINAL_ANSWER<ANSWER> where FINAL_ANSWER is your
answer. For example, if your answer is A, you should say <ANSWER>A<ANSWER>
7.4 Parabolic Trajectory
We use the same prompt for all of the puzzles.
You are given an image showing part of a ball’s parabolic trajectory (in red) as it moves through the air. The ball will eventually land on a horizontal surface shown in blue. Along this surface, four positions are marked: A, B, C, and D (from right to left). Question: Based on the shape and direction of the visible trajectory, extrapolate the rest of the trajectory and determine which point (A, B, C, or D) is the first likely impact point of the ball on the surface?
You can explain your reasoning but you should specify your final answer by putting <ANSWER> before and after it like <ANSWER>FINAL_ANSWER<ANSWER> where FINAL_ANSWER is your
answer. For example, if your answer is A, you should say <ANSWER>A<ANSWER>
8 Examples of Hallucinations in the Models
In this part, we provide examples of hallucination and fix answers that some models provide.
8.1 Seven Segments
For this puzzle, some models tend to repeatedly produce a fixed answer, such as , regardless of changes in the input. These default responses vary across models, and there is no single answer that is universally repeated. Additionally, some of the generated answers contain a different number of digits than what is specified in the prompt, indicating a failure to follow basic task constraints. Table 6 summarizes this behavior for a subset of models that exhibit these patterns.
| Model | Difficulty | Answer | Number of occurrence (out of ) | |||
| Claude 3.7 Sonnet | Easy | 409 | 20 | |||
| Medium | 2023 | 60 | ||||
| Hard | 20235 | 10 | ||||
| Qwen VL 2.5 7B | Easy | 012 | 32 | |||
| Medium | 0123 | 26 | ||||
| Hard | 01234 | 18 | ||||
| Qwen VL 2.5 32B | Easy | 888 | 29 | |||
| Medium | 2023 | 25 | ||||
| Hard | 22222 or 2023 | 17 | ||||
| LLaVA-OneVision 7B | Easy | 123 | 50 | |||
| Medium | 0123 | 25 | ||||
| Hard | 12345 | 22 | ||||
| Molmo 7B | Easy | 170 or 161 | 28 | |||
| Medium | 2300 or 2310 | 25 | ||||
| Hard | 272829 | 32 |
8.2 Connect the Dots
In this part, we provide examples of answers in which the model hallucinates non-existent features. Example 1: We showed Figure 4(a) to Molmo 7B, and in its response, it is hallucinating features of a swan:
Molmo: The image appears to show a dotted outline of a bird in flight. The shape is elongated and curved, with a distinctive neck and head form that’s characteristic of certain waterfowl. The overall silhouette suggests a graceful, soaring posture.
Given the options provided, the bird that most closely matches this dotted outline is:
<ANSWER>A<ANSWER>
This dotted shape best represents a swan in flight. Swans are known for their long, elegant necks and graceful forms, which align well with the curved outline visible in the dotted image. The elongated body and the way the neck curves upwards to form the head are particularly indicative of a swan’s silhouette.
Example 2: We showed Figure 4(b) to Llama 3.2 11B, and it hallucinated onion features:
Llama: To determine the correct answer, let’s analyze the shape formed by the dots. The dots outline a teardrop shape with a narrow top and a wider bottom. This shape is most similar to an onion, which has a similar teardrop shape with a narrow top and a wider bottom.
The other options do not match the shape formed by the dots. A carrot is typically long and thin, an onion is bulbous with a narrow top and a wider bottom, and a pineapple is round with a crown on top.
Therefore, based on the shape formed by the dots, the correct answer is:
<ANSWER>B<ANSWER>"
Example 3: We showed Figure 4(a) to Claude 3.7 Sonnet, and it hallucinated features of a swan:
Claude: Looking at the dotted outline in the image, I can see a shape that appears to have:
- A long, curved neck
- A rounded body
- What looks like a tail at the back
- A distinctive shape at the bottom that could represent feet or a base
The overall silhouette strongly resembles a bird with a long, curved neck. Among the options:
A: Swan - Swans have long, curved necks and a distinctive body shape that matches this outline quite well
@@ 8.3 Extrapolation Puzzles
9 Ablation on Prompting
In this section, we explore different prompting techniques and their effects on models’ performance on Hyperphantasia.
9.1 Investigating the prompt sensitivity of the models
To ensure fairness in comparison, we conduct an experiment with different prompt formatting to investigate the effects of prompt formatting on models’ performance. For this purpose, we curated four alternative prompts and used them for answering the Easy Linear Trajectory task. Table 7 provides the curated alternative prompts, and the results are in Table 8.
| # | Change | Prompt |
| 0 | Original |
The image shows a triangle composed of two blue edges and one red edge. A portion of the red edge has been removed. The red line originates from the left red vertex and should connect to a point along the blue edge on the right side of the triangle. There are four possible points marked on this blue edge: A, B, C, and D. Question: Based on the direction and alignment of the red edge, extrapolate the rest of the red line and determine which point (A, B, C, or D) is the vertex of the triangle corresponding to the missing part of the red edge? Select the most geometrically consistent option.
You can explain your reasoning but you should specify your final answer by putting <ANSWER> before and after it like <ANSWER>FINAL_ANSWER<ANSWER> where FINAL_ANSWER is your answer. For example, if your answer is A, you should say <ANSWER>A<ANSWER. |
| 1 | Shorten |
Based on the incomplete edge of the triangle in the image, which point among A, B, C, and D is a vertex of the triangle?
You can explain your reasoning but you should specify your final answer by putting <ANSWER> before and after it like <ANSWER>FINAL_ANSWER<ANSWER> where FINAL_ANSWER is your answer. For example, if your answer is A, you should say <ANSWER>A<ANSWER. |
| 2 | Instruct the model for extra care and attention |
Carefully and skillfully review the image and answer the following question.
The image shows a triangle composed of two blue edges and one red edge. A portion of the red edge has been removed. The red line originates from the left red vertex and should connect to a point along the blue edge on the right side of the triangle. There are four possible points marked on this blue edge: A, B, C, and D. Question: Based on the direction and alignment of the red edge, extrapolate the rest of the red line and determine which point (A, B, C, or D) is the vertex of the triangle corresponding to the missing part of the red edge? Carefully, select the most geometrically consistent option. You can explain your reasoning but you should specify your final answer by putting <ANSWER> before and after it like <ANSWER>FINAL_ANSWER<ANSWER> where FINAL_ANSWER is your answer. For example, if your answer is A, you should say <ANSWER>A<ANSWER. |
| 3 | Instruct the model for single letter answering |
The image shows a triangle composed of two blue edges and one red edge. A portion of the red edge has been removed. The red line originates from the left red vertex and should connect to a point along the blue edge on the right side of the triangle. There are four possible points marked on this blue edge: A, B, C, and D. Question: Based on the direction and alignment of the red edge, extrapolate the rest of the red line and determine which point (A, B, C, or D) is the vertex of the triangle corresponding to the missing part of the red edge? Select the most geometrically consistent option.
Your answer should be a single letter: A, B, C, or D without any explanation or additional text. |
| 4 | Define AI expert |
You are an expert mental visualization AI with deep visual thinking capabilities.
The image shows a triangle composed of two blue edges and one red edge. A portion of the red edge has been removed. The red line originates from the left red vertex and should connect to a point along the blue edge on the right side of the triangle. There are four possible points marked on this blue edge: A, B, C, and D. Question: Based on the direction and alignment of the red edge, extrapolate the rest of the red line and determine which point (A, B, C, or D) is the vertex of the triangle corresponding to the missing part of the red edge? Select the most geometrically consistent option. You can explain your reasoning but you should specify your final answer by putting <ANSWER> before and after it like <ANSWER>FINAL_ANSWER<ANSWER> where FINAL_ANSWER is your answer. For example, if your answer is A, you should say <ANSWER>A<ANSWER. |
| Model | Prompt 0 | Prompt 1 | Prompt 2 | Prompt 3 | Prompt 4 | |
| Qwen VL 2.5 7B | 27 | 32 | 26 | 23 | 25 | |
| Qwen VL 2.5 32B | 40 | 38 | 35 | 35 | 27 | |
| Llama 3.2 11B | 36 | 43 | 39 | 27 | 41 | |
| Llama 3.2 90B | 30 | 28 | 32 | 21 | 27 | |
| LLaVA-OneVision 7B | 26 | 40 | 32 | 41 | 24 | |
| LLaVA-OneVision 72B | 42 | 43 | 26 | 39 | 34 |
These results highlight the fact that the prompt formatting has little effect on the performance of the models and it cannot solve the underlying problem.
9.2 Multi-shot prompting
To further explore the effects of prompting, we conducted an additional experiment with 3-shot prompting. Similar to the previous setup, we use the Easy Linear Trajectory puzzles for this experiment. Table 9 shows the results of this experiment. We do not report the 3-shot accuracy of LLaVA-OneVision 72B, as it was unable to produce a meaningful answer when multiple images were included in the prompt. We speculate that this may be due to the lack of multi-image inputs in its training data. Interestingly, this issue does not arise with the smaller version of the model, which was able to provide valid answers.
| Model | Zero-Shot | 3-Shot | |
| o4-mini | 43 | 40 | |
| Qwen VL 2.5 7B | 27 | 30 | |
| Qwen VL 2.5 32B | 40 | 31 | |
| LLaVA-OneVision 7B | 26 | 29 | |
| LLaVA-OneVision 72B | 42 | - | |
| Llama 3.2 11B | 36 | 35 | |
| Llama 3.2 90B | 30 | 31 |
As with the previous experiment, we observe minimal to no improvement in model performance, suggesting that multi-shot prompting alone is insufficient to overcome the fundamental limitations of current models in mental visualization tasks.
10 Language-only Seven Segments task
Seven Segments puzzles are the only task in Hyperphantasia that can be completely explained in a language-only format without showing the image. As a result, we curate a new text-only prompt for the Easy Seven Segments puzzles and use it to test some of the previously tested models and language-only models. Here is the curated prompt:
Solve the following puzzle:
Imagine a 3 by 6 grid of dots arranged as follows (rows: top to bottom, columns: left to right):
Column 1: [0] [1] [2]
Column 2: [3] [4] [5]
Column 3: [6] [7] [8]
Column 4: [9] [10] [11]
Column 5: [12] [13] [14]
Column 6: [15] [16] [17]
Now, imagine the following pairs of dots are connected with straight lines:
<CONNECTIONS>
What 3-digit number is formed by these connections?
You can explain your reasoning but you should specify your final answer by putting <ANSWER> before and after it like <ANSWER>FINAL_ANSWER<ANSWER> where FINAL_ANSWER is your answer.
Your answer should be a three digit number formed by the connections.
In the prompt, <CONNECTIONS> refers to the specific connections of each puzzle. Table 10 summarizes the results of this experiment.
| Model | Model type | Accuracy in text-only format | Accuracy in image format | |
| o4-mini | VL | 88 | 83 | |
| GPT-4o | VL | 1 | 4 | |
| Qwen VL 2.5 7B | VL | 0 | 0 | |
| Qwen VL 2.5 32B | VL | 1 | 0 | |
| Qwen 2.5 32B Instruct | L | 0 | - | |
| DeepSeek-R1 distill Qwen 32B | L | 0 | - | |
| Llama 3 8B Instruct | L | 0 | - |
These results demonstrate that language-only models are unable to solve this task. This outcome is expected, as these models are not trained on any visual data and therefore lack the mental visualization capabilities. Additionally, we observe that the multi-modal models which previously succeeded on this task (such as o4-mini and, to a lesser extent, GPT-4o) continue to solve it even in the text-only format. Meanwhile, other models that initially failed still struggle with this version of the task.